# Self-Ensembling Gaussian Splatting for Few-shot Novel View Synthesis

### v1

@@ Abstract
3D Gaussian Splatting (3DGS) has demonstrated remarkable effectiveness for novel view synthesis (NVS). However, the 3DGS model tends to overfit when trained with sparse posed views, limiting its generalization capacity for broader pose variations. In this paper, we alleviate the overfitting problem by introducing a self-ensembling Gaussian Splatting (SE-GS) approach. We present two Gaussian Splatting models named the $\mathbf{\Sigma}$-model and the $\mathbf{\Delta}$-model. The $\mathbf{\Sigma}$-model serves as the primary model that generates novel-view images during inference. At the training stage, the $\mathbf{\Sigma}$-model is guided away from specific local optima by an uncertainty-aware perturbing strategy. We dynamically perturb the $\mathbf{\Delta}$-model based on the uncertainties of novel-view renderings across different training steps, resulting in diverse temporal models sampled from the Gaussian parameter space without additional training costs. The geometry of the $\mathbf{\Sigma}$-model is regularized by penalizing discrepancies between the $\mathbf{\Sigma}$-model and the temporal samples. Therefore, our SE-GS conducts an effective and efficient regularization across a large number of Gaussian Splatting models, resulting in a robust ensemble, the $\mathbf{\Sigma}$-model. Experimental results on the LLFF, Mip-NeRF360, DTU, and MVImgNet datasets show that our approach improves NVS quality with few-shot training views, outperforming existing state-of-the-art methods. The code is released at:https://github.com/sailor-z/SE-GS.
@@ 1 Introduction
Novel view synthesis (NVS) is a critical task [49] in computer vision and graphics, playing a pivotal role in applications such as virtual reality [8], augmented reality [50], and 3D content generation [15, 42]. The objective of NVS is to generate photo-realistic images from previously unseen viewpoints. Typically, NVS starts by constructing a 3D representation [26, 39] from a set of existing 2D observations. In recent years, 3D Gaussian Splatting (3DGS) [20, 6, 47] has emerged as a powerful representation, integrating the advantages of both explicit [33] and implicit [26] representations. This approach enables efficient novel view generation and yields promising synthesized results with densely sampled observations that cover a wide range of viewpoints.
@@
Figure 2: Overfitting in 3D Gaussian Splatting with few-shot training views. 2(a) and 2(b) illustrate the performance of 3DGS on training and testing views, respectively. Each curve represents the PSNR values across different training iterations.
However, 3DGS tends to overfit the available views when only a limited number of images are provided. As shown in Fig. 2, we evaluate the 3DGS model trained on sparse images with different numbers of iterations. The performance on the training data consistently improves as the number of iterations increases, while the testing results deteriorate after 2000 iterations. Moreover, the overfitting issue becomes more noticeable with fewer training views, such as when using only 3 views. To address this problem, we introduce a new 3DGS approach in this paper, which enhances the quality of novel view synthesis with sparse training views via self-ensembling. Self-ensembling [11, 21, 27] has been witnessed in the literature as an effective strategy to mitigate overfitting, particularly in scenarios where limited data is available [45]. This technique leverages multiple predictions from the model under different conditions such as dropout [37] and data augmentation [36], forming a consensus prediction via regularization. By aggregating outputs with diverse perturbations, self-ensembling creates a more stable target for training, reducing the noise in individual predictions. Despite its success, how to utilize self-ensembling in the 3DGS pipeline remains an open problem. Therefore, we explore this gap, introducing a new 3DGS training pipeline consisting of a $\mathbf{\Sigma}$-model and a $\mathbf{\Delta}$-model.
@@
Specifically, we present a $\mathbf{\Delta}$-model that represents a temporal sample in the Gaussian parameter space. Notably, in the scenario of sparse views, the training data lacks sufficient information to identify an optimal sample in the parameter space. As a result, the training process tends to converge to a local optimum, leading to an overfitting problem. To tackle this issue, one could create more samples by training multiple $\mathbf{\Delta}$-models, thereby increasing the likelihood of reaching a global optimum. Nevertheless, as we will demonstrate in Sec. 4, this approach incurs significant computational costs, limiting its ability to produce diverse samples. In contrast, we train a single $\mathbf{\Delta}$-model following the typical 3DGS training strategy [20], while obtaining various samples via perturbing the $\mathbf{\Delta}$-model. Rather than applying random perturbations, which often result in invalid samples, we perturb the $\mathbf{\Delta}$-model based on uncertainties derived from the training data. We store images generated by rendering the $\mathbf{\Delta}$-model from a set of pseudo views in buffers at different training iterations. These pseudo views are randomly sampled from the camera trajectories of the training views. We assess the reliability of the renderings by calculating the pixel-level uncertainties across the images within each buffer. Pixels with large uncertainty scores are identified, and the Gaussian parameters of the $\mathbf{\Delta}$-model associated with these pixels are perturbed by random noise. Therefore, our perturbation strategy generates uncertainty-aware temporal samples without introducing significant training overhead.
@@
Building upon the $\mathbf{\Delta}$-model, we construct an ensemble of the temporal samples by proposing a $\mathbf{\Sigma}$-model. We train the $\mathbf{\Sigma}$-model over the training views without the aforementioned perturbations. To achieve self-ensembling, we regularize the geometry of the $\mathbf{\Sigma}$-model, minimizing the discrepancies between the $\mathbf{\Sigma}$-model and the perturbed $\mathbf{\Delta}$-model. The discrepancies are measured through a photometric loss between images synthesized from the pseudo views. During testing, we employ the $\mathbf{\Sigma}$-model as the primary model for novel view synthesis. Since the regularization is performed based on diverse samples in the Gaussian parameter space, the self-ensembling process enhances the robustness of the $\mathbf{\Sigma}$-model, thereby improving the generalization ability to novel views.
@@
We conduct experiments on several datasets consisting of LLFF [25], DTU [19], Mip-NeRF360 [3], and MVImgNet [46]. The evaluation is performed with varying numbers of training views. Our SE-GS achieves the best performance across all the datasets when limited training views are used, surpassing the state-of-the-art approaches. Additionally, we provide a comprehensive analysis in which the results demonstrate the effectiveness and efficiency of our method. To the best of our knowledge, we are the first to explore the potential of the ensembling mechanism in 3DGS for few-shot novel view synthesis. We present a new self-ensembling technique that effectively mitigates the overfitting of 3DGS without requiring additional ground-truth signals or incurring significant training overhead.
@@ 2 Related Work
Radiance field modeling. Recently, Neural Radiance Fields (NeRF) [26, 41, 30, 3] has brought groundbreaking advancements to novel view synthesis. NeRF learns the density and color information of a scene or object from multi-view posed images, enabling the synthesis of realistic images from novel views via volume rendering [9]. However, most NeRF-based methods are inefficient [17] during training and inference because they require numerous MLP queries in the rendering process. This limitation makes it challenging to support tasks with real-time requirements. To this end, Kebr et al. [20] proposes an efficient alternative for radiance field modeling, which enables neural rendering via 3D Gaussian Spaltting (3DGS). 3DGS relies on explicit representations that allow for faster rendering and training [10]. Some methods [40, 5, 24, 7] explore the generalization ability of 3DGS towards novel scenes, optimizing the Gaussian parameters based on learnable cost volumes [14]. In this paper, we stick to the vanilla 3DGS rendering pipeline, focusing on the scene-level radiance field modeling.
@@
3DGS with few-shot images. The inadequacy of training views is a primary factor leading to artifacts in novel view synthesis. Even when training views are relatively dense, some regions may still be observed from few-shot images, resulting in quality degradation in the synthesized views. To handle such an under-constrained problem, some approaches incorporate additional ground-truth signals into the 3DGS pipeline. For instance, DNGaussian [23], CoherentGS [29], and FSGS [53] utilize a monocular depth estimator [32] to predict ground-truth depth maps, enabling the 3DGS model to be trained with both photometric and depth loss. Other methods, such as [43] and [16], leverage multi-view stereo techniques [35] to generate novel-view images as ground truth. However, ground-truth data obtained from off-the-shelf methods is inevitably noisy, which potentially impacts the training of 3DGS. As noted in [48] and observed in our experiments, when the number of training views increases, the performance of these methods is sometimes worse than the vanilla 3DGS. In this context, Zhang et al. [48] proposes training multiple 3DGS models with a cross-model regularization term. This method mitigates the overfitting problem without relying on additional data. Nevertheless, training additional 3DGS models leads to significant computational costs, limiting the effectiveness of the regularization.
@@
Ensembling. Ensembling has been witnessed as a powerful technique [52, 1, 13] in machine learning to improve model robustness and generalization by aggregating predictions from multiple models. Traditionally, ensembles [12, 31, 22] are created by training independent models and averaging their outputs or applying a voting mechanism. To improve the efficiency, some self-ensembling methods, such as temporal ensembling [21] and consistency regularization [11], leverage variations of a single model across training iterations to build an ensemble. Motivated by the success of self-ensembling, we present a Self-Ensembling Gaussian Splatting approach, in which we generate diverse samples from the Gaussian parameter space through an uncertainty-aware perturbation mechanism.
@@ 3 Method
Figure 3: Pipeline of the presented SE-GS. We tackle the scenario where only sparse posed images are available. To mitigate the overfitting of 3DGS, we incorporate a self-ensembling mechanism into the 3DGS framework. We present a $\mathbf{\Delta}$-model that represents a temporal sample in the Gaussian parameter space. This $\mathbf{\Delta}$-model is perturbed during training, which results in diverse samples. The perturbation process is carried out with the guidance of uncertainties computed from images rendered from pseudo views. The pixels in the image buffers with high uncertainties are identified and the Gaussians overlapping these pixels are perturbed, highlighted as red ellipses. The self-ensembling over the temporal samples is achieved by training a $\mathbf{\Sigma}$-model with a regularization that penalizes the discrepancies of the $\mathbf{\Sigma}$-model and the perturbed $\mathbf{\Delta}$-model. During inference, the novel view synthesis is performed using the $\mathbf{\Sigma}$-model.
@@ 3.1 Preliminaries and Motivations
3D Gaussian Splatting [20] represents a scene as a collection of Gaussians. These Gaussians are splatted onto the 2D image plane during rendering. Formally, we denote the set of $N$ Gaussians as $\{{G_{i}},i=1,2,...,N\}$. Each Gaussian $G_{i}$ is defined by $G_{i}=(\mathbf{\mu}_{i},\Sigma_{i},\mathbf{h}_{i},o_{i})$, where $\mathbf{\mu}_{i}$ is the center of the Gaussian in 3D space, $\Sigma_{i}$ is the covariance matrix, $\mathbf{h}_{i}$ represents spherical harmonics (SH) coefficients associated with the Gaussian, and $o_{i}$ indicates the opacity. Each Gaussian contributes to a 3D point $\mathbf{x}$ according to a 3D Gaussian distribution:
$G_{i}(\mathbf{x})=e^{-\frac{1}{2}(\mathbf{x}-\mathbf{\mu}_{i})^{T}\Sigma_{i}^{-1}(\mathbf{x}-\mathbf{\mu}_{i})}.$ (1)
To ensure that $\Sigma_{i}$ remains positive semi-definite throughout optimization, this matrix is decomposed into two learnable components as $\Sigma_{i}=\mathbf{R}\mathbf{S}\mathbf{S}^{T}\mathbf{R}^{T}$, where $\mathbf{R}$ is a rotation matrix and $\mathbf{S}$ stands for a scaling matrix. For RGB rendering, the projection from 3D space to a 2D image plane is defined as:
@@
where $\Sigma_{i}^{{}^{\prime}}$ is the projected 2D covariance matrix, $\mathbf{W}$ represents the projection matrix, and $\mathbf{J}$ denotes the Jacobian of the affine approximation of the projection. The color of a pixel is obtained by performing alpha-blending as
@@
where $M$ is the number of Gaussians covering the pixel, $\mathbf{c}_{i}$ denotes the color of the Gaussian derived from the SH coefficients, and $\alpha_{i}$ is computed from the 2D covariance matrices and opacity scores. For better convergence, the Gaussian parameters are initialized based on the 3D points obtained by employing structure-from-motion (SFM) techniques [33, 34]. These parameters are optimized via gradient descent over the available posed training images.
In this paper, we focus on the scenario of sparse views, where only few-shot training images are known. We denote the 3DGS model trained on these images as $\mathcal{G}$, which represents a sample in the Gaussian parameter space. In this setting, $\mathcal{G}$ is prone to overfitting the training data and thereby getting trapped in local optima, which ultimately degrades the quality of the synthesized novel views.
@@
To handle the overfitting problem, a promising solution [48] is to train a set of 3DGS models $\{\mathcal{G}_{1},\mathcal{G}_{2},...,\mathcal{G}_{k}\}$ that correspond to different samples. By applying regularization across this set, each model is guided away from specific local optima, facilitating a more thorough exploration of the parameter space. As we will report in Sec. 4, the performance on testing data improves as $k$ grows. However, training multiple 3DGS models incurs significant computational costs, which become unaffordable with a large $k$. This limitation therefore acts as an obstacle to scaling up the method for better performance. Drawing inspiration from the success of the regularization, we propose a novel training strategy based on self-ensembling [11, 21], which enhances 3DGS for few-shot NVS with negligible additional training overhead. The pipeline is illustrated in Fig. 3 and the details will be elaborated in this section.
3.2 Temporal Samples in 3DGS

### v2

@@ Abstract
3D Gaussian Splatting (3DGS) has demonstrated remarkable effectiveness for novel view synthesis (NVS). However, the 3DGS model tends to overfit when trained with sparse posed views, limiting its generalization ability to novel views. In this paper, we alleviate the overfitting problem, presenting a Self-Ensembling Gaussian Splatting (SE-GS) approach. Our method encompasses a $\mathbf{\Sigma}$-model and a $\mathbf{\Delta}$-model. The $\mathbf{\Sigma}$-model serves as an ensemble of 3DGS models that generates novel-view images during inference. We achieve the self-ensembling by introducing an uncertainty-aware perturbation strategy at the training state. We complement the $\mathbf{\Sigma}$-model with the $\mathbf{\Delta}$-model, which is dynamically perturbed based on the uncertainties of novel-view renderings across different training steps. The perturbation yields diverse temporal samples in the Gaussian parameter space without additional training costs. The geometry of the $\mathbf{\Sigma}$-model is regularized by penalizing discrepancies between the $\mathbf{\Sigma}$-model and these temporal samples. Therefore, our SE-GS conducts an effective and efficient regularization across a large number of 3DGS models, resulting in a robust ensemble, the $\mathbf{\Sigma}$-model. Our experimental results on the LLFF, Mip-NeRF360, DTU, and MVImgNet datasets show that our approach improves NVS quality with few-shot training views, outperforming existing state-of-the-art methods. The code is released at:project page.
@@ 1 Introduction
Novel view synthesis (NVS) is a critical task [49] in computer vision and graphics, playing a pivotal role in applications such as virtual reality [8], augmented reality [50], and 3D content generation [16, 42]. The objective of NVS is to generate photo-realistic images from previously unseen viewpoints. Typically, NVS starts by constructing a 3D representation [28, 39] from a set of existing 2D observations. In recent years, 3D Gaussian Splatting (3DGS) [21, 6, 46] has emerged as a powerful representation, integrating the advantages of both explicit [35] and implicit [28] representations. This approach enables efficient novel view generation and yields promising synthesized results with densely sampled observations that cover a wide range of viewpoints. However, 3DGS tends to overfit the available views when only a limited number of images are provided. To illustrate this issue, in Fig. 2, we evaluate the 3DGS model trained on sparse images with different numbers of iterations. The performance on the training data consistently improves as the number of iterations increases, while the testing results deteriorate after 2000 iterations. Moreover, the overfitting problem becomes more noticeable with fewer training views, such as when using only 3 views.
@@
Figure 2: Overfitting in 3D Gaussian Splatting with sparse training views. 2(a) and 2(b) illustrate the performance of 3DGS on training and testing views, respectively. Each curve represents the PSNR values across different training iterations.
@@
To mitigate overfitting, self-ensembling [11, 22, 29] has been highlighted as an effective strategy in tasks such as detection [48] and segmentation [25]. Nevertheless, its use for NVS remains unstudied, and how to exploit it in a 3DGS formalism is an open question. Therefore, in this paper, we bridge this gap, introducing a new 3DGS method that enhances the quality of novel view synthesis with sparse training views via self-ensembling. Our method consists of two models, a $\mathbf{\Sigma}$-model and a $\mathbf{\Delta}$-model. The $\mathbf{\Sigma}$-model constitutes an ensemble of 3DGS models, which is used to generate novel views during inference. A straightforward way to achieve ensembling would be to utilize multiple 3DGS models that correspond to different samples in the Gaussian parameter space, thereby increasing the likelihood of reaching a more robust solution. Nevertheless, as we will demonstrate in Sec. 4, the computational cost of training multiple 3DGS models limits the ability to produce sufficiently diverse samples. By contrast, we complement the $\mathbf{\Sigma}$-model with a single $\mathbf{\Delta}$-model and obtain diverse samples by perturbing the $\mathbf{\Delta}$-model. Specifically, instead of applying random perturbation, which often results in invalid samples, we perturb the $\mathbf{\Delta}$-model based on uncertainties derived from the training data. We store images generated by rendering the $\mathbf{\Delta}$-model from a set of pseudo views at different training iterations in buffers and calculate pixel-level uncertainties across images within each buffer. We then add random noise to the Gaussian parameters of the $\mathbf{\Delta}$-model associated with pixels that have high uncertainty scores. Consequently, our perturbation strategy generates uncertainty-aware samples without introducing significant training overhead.
@@
Building upon the uncertainty-aware samples derived from the $\mathbf{\Delta}$-model, we construct an ensemble by leveraging a self-ensembling mechanism. We train the $\mathbf{\Sigma}$-model over the training views without the aforementioned perturbation and regularize the geometry of the $\mathbf{\Sigma}$-model by minimizing its discrepancies with the perturbed $\mathbf{\Delta}$-model. These discrepancies are measured via a photometric loss between images synthesized from the pseudo views. Since the regularization is performed based on diverse samples in the Gaussian parameter space, the self-ensembling process enhances the robustness of the $\mathbf{\Sigma}$-model, thereby improving its generalization ability to novel views. Moreover, our self-ensembling technique is independent of any additional ground-truth signals as the regularization is carried out in a self-supervised manner.
@@
We conduct experiments on multiple datasets, including LLFF [27], DTU [20], Mip-NeRF360 [3], and MVImgNet [45], with limited numbers of training views. Our SE-GS achieves the best performance across all datasets in the sparse-view setting, surpassing the state-of-the-art approaches. Furthermore, we perform a comprehensive analysis of our method to demonstrate its effectiveness and efficiency. To the best of our knowledge, we are the first to explore the potential of the self-ensembling mechanism in 3DGS for few-shot novel view synthesis. The code will be publicly available upon acceptance.
@@ 2 Related Work
Radiance field modeling. Recently, Neural Radiance Field (NeRF) [28, 41, 32, 3] has brought groundbreaking advancements to novel view synthesis. NeRF learns the density and color information of a scene or object from multi-view posed images, enabling the synthesis of realistic images from novel views via volume rendering [9]. However, most NeRF-based methods are inefficient [18] during training and inference because they require numerous MLP queries in the rendering process. This limitation makes it challenging to support tasks with real-time requirements. In this context, 3D Gaussian Splatting [21] is developed as an efficient alternative for radiance field modeling. 3DGS relies on explicit representations [35] that allow for faster rendering and training [10]. While initially introduced as a scene-specific model, some 3DGS variants [40, 5, 26, 7] explore the generalization ability of 3DGS towards novel scenes, optimizing the Gaussian parameters based on learnable cost volumes [15]. In this paper, we stick to the vanilla 3DGS rendering pipeline, focusing on the scene-level radiance field modeling.
@@
3DGS with few-shot images. The insufficiency of training views is a primary factor leading to artifacts in novel view synthesis. Even when training views are relatively dense, some regions may still be observed from few-shot images, resulting in quality degradation in the synthesized views. To handle such an under-constrained problem, some approaches incorporate additional ground-truth signals into the 3DGS pipeline. For instance, DNGaussian [24], CoherentGS [31], and FSGS [53] utilize a monocular depth estimator [34] to predict depth maps, enabling the 3DGS model to be trained with both a photometric and a depth loss. Other methods, such as [43] and [17], leverage multi-view stereo techniques [37] to generate novel-view images as ground truth. However, ground-truth data obtained from off-the-shelf methods is inevitably noisy, which potentially impacts the training of 3DGS. As noted in [47] and observed in our experiments, when the number of training views increases, the performance of these methods is sometimes worse than the vanilla 3DGS. To overcome this problem, Zhang et al. [47] propose training multiple 3DGS models with a cross-model regularization term. However, training additional 3DGS models incurs a significant computational cost, making it impractical to scale up to a large number of 3DGS models for stronger regularization.
@@
Ensembling. Ensembling has been evidenced as a powerful technique in machine learning [52, 1, 14] to improve model robustness and generalization by aggregating predictions from multiple models. Traditionally, ensembles [13, 33, 23] are created by training independent models and averaging their outputs or applying a voting mechanism. To improve the efficiency, some self-ensembling methods, such as temporal ensembling [22] and consistency regularization [11], leverage variations of a single model across training iterations to build an ensemble. Motivated by the success of self-ensembling, we present the first Self-Ensembling Gaussian Splatting approach, in which we generate diverse samples in the Gaussian parameter space through an uncertainty-aware perturbation mechanism.
@@ 3 Method
Figure 3: Pipeline of the presented SE-GS. We tackle the overfitting problem in sparse-view scenarios by incorporating a self-ensembling mechanism into the 3DGS framework. Our self-ensembling builds upon diverse samples in the Gaussian parameter space. These samples are derived from the $\mathbf{\Delta}$-model via an uncertainty-aware perturbation strategy. Specifically, we store images rendered from pseudo views at different training steps in buffers, from which we compute pixel-level uncertainties. We then perturb the Gaussians overlapping the pixels with high uncertainties, as highlighted as red ellipses. Self-ensembling over the resulting samples is achieved by training a $\mathbf{\Sigma}$-model with a regularization that penalizes the discrepancies of the $\mathbf{\Sigma}$-model and the perturbed $\mathbf{\Delta}$-model. During inference, novel view synthesis is performed using the $\mathbf{\Sigma}$-model.
@@ 3.1 Preliminaries and Motivations
3D Gaussian Splatting [21] represents a scene as a collection of Gaussians. These Gaussians are splatted onto the 2D image plane during rendering. Formally, we denote the set of $N$ Gaussians as $\{{G_{i}},i=1,2,...,N\}$. Each Gaussian $G_{i}$ is defined as $G_{i}=(\mathbf{\mu}_{i},\Sigma_{i},\mathbf{h}_{i},o_{i})$, where $\mathbf{\mu}_{i}$ is the 3D position, $\Sigma_{i}$ is the covariance matrix, $\mathbf{h}_{i}$ represents spherical harmonics (SH) coefficients associated with the Gaussian, and $o_{i}$ indicates opacity. Each Gaussian contributes to a 3D point $\mathbf{x}$ according to the 3D Gaussian distribution
$\mathcal{N}_{3d}^{i}(\mathbf{x})=e^{-\frac{1}{2}(\mathbf{x}-\mathbf{\mu}_{i})^{T}\Sigma_{i}^{-1}(\mathbf{x}-\mathbf{\mu}_{i})}.$ (1)
The parameters of the Gaussians are then learned from the input views. To ensure that $\Sigma_{i}$ remains positive semi-definite throughout optimization, it is decomposed into two learnable components as $\Sigma_{i}=\mathbf{R}\mathbf{S}\mathbf{S}^{T}\mathbf{R}^{T}$, where $\mathbf{R}$ is a rotation matrix and $\mathbf{S}$ stands for a scaling matrix. For RGB rendering, the projection from 3D space to a 2D image plane relies on the projection matrix $\mathbf{W}$ to compute the projected 2D covariance matrix
@@
where $\mathbf{J}$ denotes the Jacobian of the affine approximation of the projection. The color of a pixel is obtained by performing alpha-blending as
@@
where $M$ is the number of Gaussians covering the pixel, $\mathbf{c}_{i}$ denotes the color of the Gaussian derived from the SH coefficients, and $\alpha_{i}$ is computed from the 2D covariance matrices and opacity scores. For better convergence, the Gaussian parameters are initialized based on 3D points obtained using structure-from-motion (SFM) techniques [35, 36]. These parameters are optimized via gradient descent with a photometric loss [21] where the posed training images serve as ground truth.
@@
In this paper, we focus on sparse-view scenarios, where only few-shot training images are provided. We denote the 3DGS model trained on these images as $\mathcal{G}$, which represents a sample in the Gaussian parameter space. In this setting, $\mathcal{G}$ is prone to overfitting the training data and thereby getting trapped in a suboptimal solution, which ultimately degrades the quality of the synthesized novel views. To handle the overfitting problem, a promising solution is to train a set of 3DGS models $\{\mathcal{G}_{1},\mathcal{G}_{2},...,\mathcal{G}_{k}\}$ that correspond to different samples. By applying regularization across this set, each model is guided away from specific local optima, facilitating a more thorough exploration of the parameter space. As we will report in Sec. 4, the performance on testing data improves as $k$ grows. However, training multiple 3DGS models incurs a significant computational cost, which becomes unaffordable with a large $k$. This limitation prevents the method from scaling up for better performance. Consequently, we propose a novel training strategy based on self-ensembling [11, 22], which enhances 3DGS for few-shot NVS with negligible additional training overhead. The pipeline is illustrated in Fig. 3 and the details will be elaborated in the remainder of this section.
3.2 Uncertainty-Aware Perturbation